Ask
your own question, for FREE!
Mathematics
16 Online
if two dice are rolled what is the probability of getting sum of 6 given that atleast one dice shows 2?
Still Need Help?
Join the QuestionCove community and study together with friends!
Probability the dice shows 2=1/6 If at least one dice show two and getting the sum 6 then it will be either 4,2 or 2,4 so the answer is 2(1/6)(1/6)
didnt get it two dice are rolled.
now the condition: 1. Sum=6 2. At least one must be 2 so there is two possibility: 1. we get 2 first and then get 4 2.we get 4 first and then get 2 Probability of 1.: (1/6)(1/6)=1/36 Probability of 2.: (1/6)(1/6)=1/36 Total probability= 1/36+1/36=1/18
i'm not too good at probability but isn't this conditional probability? you are given that 1 dice is 2 so should not the 1/6 's be added?
@cwrw238 you are right.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
15 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal